OpenStudy (anonymous):

what is the meaning of |x| ≤ 0?

OpenStudy (saifoo.khan):

\[x \le 0, x \ge0\]

OpenStudy (anonymous):

It means the distance on the number line between the point x and 0 is less than or equal to 0.

OpenStudy (anonymous):

so x=0

OpenStudy (anonymous):

how do i graph tht on a number line?

OpenStudy (anonymous):

If \(|x| \le 0\), then \(x = 0\) The absolute value cannot be negative

OpenStudy (anonymous):

distance cannot be negative. it's just the point 0.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

no numbr line

OpenStudy (anonymous):



OpenStudy (anonymous):

thnk u so much

OpenStudy (anonymous):

what polpak drew was the number line :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The x with the two lines on either side means absolute value which means whatever the number is (negative or positive), you take it as a positive. The absolute value of -5 is 5 and the absolute value of 10 is 10. In your case, since the absolute value of x <= 0, then there is only one possible answer which is x = 0. All negative values would become positives and none of them are equal to or less than zero. And all positive values are not less than or equal to 0.

OpenStudy (anonymous):

y=-x+3; A(6,3)? is the ans 3=-3?